# Confidence-driven Sampling for Backdoor Attacks

### v1

@@ 4.3 Theoretical understandings
To better understand CBS, we conduct theoretical analysis on a simple SVM model. As shown in Figure 3, we consider a binary classification task where two classes are uniformly distributed in two balls centered at $\mu_{1}$(orange circle) and $\mu_{2}$(blue circle) with radius $r$ respectively in latent space [^3] This analysis is suitable for any neural networks whose last layer is a fully connected layer.:
$C_{1}\sim p_{1}(x)=\frac{1}{\pi r^{2}}1[\|x-\mu_{1}\|_{2}\leq r],\;\text{and }C_{2}\sim p_{2}(x)=\frac{1}{\pi r^{2}}1[\|x-\mu_{2}\|_{2}\leq r],$ (3)
@@
where let $\mu_{2}=0$ for simplicity. Assume that each class contains $n$ samples. We consider a simple attack that selects one single sample $x$ from class $C_{1}$, add a trigger to it to generate a poisoned $\tilde{x}$, and assign a label as class $C_{2}$ for it. Let $\tilde{C}_{1},\tilde{C}_{2}$ denote the poisoned data, and we can obtain a new backdoored decision boundary of SVM on the poisoned data. To study the backdoor effect of the trigger, we assume $\tilde{x}=x+\epsilon\frac{t}{\|t\|}$ where $\frac{t}{\|t\|},\epsilon$ denote the direction and strength of the trigger, respectively. To explain this design, we assume that the trigger introduces a ’feature’ to the original samples (Khaddaj et al., 2023), and this ’feature’ is closely related to the target class while nearly orthogonal to the prediction features [^4] Prediction feature here is referred to features used for prediction when no triggers invloved.. In addition, we assume $t$ is fixed for simplicity, which means this trigger is universal and we argue that this is valid because existing attacks such as BadNet (Gu et al., 2017) and Blend (Chen et al., 2017) inject the same trigger to every sample. To ensure the backdoor effect, we further assume $(\mu_{2}-\mu_{1})^{T}t\geq 0$, otherwise the poisoned sample will be even further from the target class (shown as the direction of the green dashed arrow) and lead to subtle backdoor effects. We are interested in two questions: Are boundary samples harder to detect? How do samples affect the backdoor performance?
@@
To investigate the first question, we adopt the Mahalanobis distance (Mahalanobis, 2018) between the poisoned sample $\tilde{x}$ and the target class $\tilde{C}_{2}$ as an indicator of outliers. A smaller distance means $\tilde{x}$ is less likely to be an outlier, indicating better stealthiness. For the second question, we estimate the success rate by estimating the volume (or area in 2D data) of the shifted class $C_{1}$ to the right of the backdoored decision boundary. This is because when triggers are added to every sample, the whole class will shift in the direction of $t$, shown as the orange dashed circle in Figure 3. The following theorem provides an estimation of Mahalanobis distance and success rate.
Theorem 4.1.
@@
Assume $\tilde{x}=x+\epsilon t/\|t\|_{2}:=x+a$ for some trigger $t$ and strength $\epsilon$, and assume $(\mu_{2}-\mu_{1})^{T}t\geq 0$. Mahalanobis distance between the poisoned sample $\tilde{x}$ and the $\tilde{C}_{2}$ is
$d_{M}^{2}(\tilde{x},\tilde{C}_{2})=\frac{4n}{(n+1)r^{2}}\frac{(n+1)r^{2}}{(n+1)r^{2}/\|\tilde{x}\|_{2}^{2}+4}.$ (4)
In addition, when training the poisoned data set using the vanilla clean SVM, taking small attack strength $\epsilon$, the success rate is an increasing function of
$sr(\tilde{x})=\epsilon\cos(t,\tilde{x}-\mu_{1})-\|\tilde{x}-\mu_{1}\|/2-r/2$
@@
Detailed proof can be found in Appendix A.1. Based on the theorem, a smaller $|\tilde{x}|_{2}$ results in a smaller $d_{M}^{2}$, making it less likely to be detected as an outlier. Additionally, a closer proximity between $\tilde{x}$ and $\mu_{1}$ corresponds to a higher success rate. To be more specific, we take two samples, $x_{1}$ close to the clean boundary (the grey dashed line), $x_{2}$ far from the boundary, as examples. It is obvious that $\tilde{x}_{1}$ is closer to center $\mu_{2}$, thus $\|\tilde{x}_{1}\|\leq\|\tilde{x}_{2}\|$, and then $d_{M}^{2}(\tilde{x}_{1},\tilde{C}_{2})\leq d_{M}^{2}(\tilde{x}_{2},\tilde{C}_{2})$ indicating that $\tilde{x}_{1}$ is harder to detect. On the other hand, as $\tilde{x}_{2}$ is closer to $\mu_{1}$, $sr(\tilde{x}_{2})\geq sr(\tilde{x}_{1})$, meaning boundary samples will have worse backdoor effect without defenses. These observations imply the trade-off between stealthiness and backdoor performance without defenses, and our experiments in Section 5 further illustrate that incorporating boundary samples can significantly improve the stealthiness with a slight sacrifice of success rate without defenses.
@@ 5 Experiment
In this section, we conduct experiments to validate the effectiveness of CBS, and show its ability to boost the stealthiness of various existing attacks. We evaluate CBS and baseline samplings under no-defense and various representative defenses in Section 5.25.3, and 5.4. In Section 5.5, we will provide more empirical evidence to illustrate that CBS is harder to detect and mitigate.
5.1 Experimental settings
To comprehensively evaluate CBS and show its ability to be applied to various kinds of attacks, we consider 3 types  [^5] We determine the types based on the threat models of attacking methods. of attacking methods that cover most of existing backdoor attacks.
In detail, Type I backdoor attacks allow attackers to inject triggers into a proportion of training data and release the poisoned data to the public. Victims train models on them from scratch. The attack’s goal is to misclassify samples with triggers as the pre-specified target class (also known as the all-to-one scenario). Type II backdoor attacks are similar to Type I and the difference is that victims finetune pre-trained models on poisoned data and the adversary’s goal is to misclassify samples from one specific class with triggers as the pre-specified target class (also known as the one-to-one scenario). Type III backdoor attacks are slightly different, and allow attackers to optimize the triggers and model parameters at the same time under the all-to-one scenario.
Baselines for sampling. We compare CBS with two baselines—Random and FUS (Xia et al., 2022). The former selects samples to be poisoned with a uniform distribution, and the latter selects samples that contribute more to the backdoor injection via computing the forgetting events (Toneva et al., 2018) for each sample. In our evaluation, we focus on image classification tasks on datasets Cifar10 and Cifar100 (Krizhevsky et al., 2009), and model architectures ResNet18 (He et al., 2016), VGG16 (Simonyan & Zisserman, 2014). We use ResNet18 as the surrogate model for CBS and FUS if not specified. The surrogate model is trained on the clean training set via SGD for 60 epochs, initial learning rate 0.01 and reduced by 0.1 after 30 and 50 epochs. We implement CBS according to Algorithm.1 and follow the original setting in (Xia et al., 2022) to implement FUS, i.e., 10 overall iterations and 60 epochs for updating the surrogate model in each iteration.
5.2 Performance of CBS in Type I backdoor attacks
Attacks & Defenses. We consider 3 representative attacks in this category—BadNet (Gu et al., 2017) which attaches a small patch pattern as the trigger to samples to inject backdoors into neural networks; Blend (Chen et al., 2017) which applies the image blending to interpolate the trigger with samples; and Adaptive backdoor [^6] Both Adaptive-Blend and Adaptive-Patch are included (Qi et al., 2022) which introduces regularization samples to improve the stealthiness of backdoors, as backbone attacks. We include 4 representative defenses: Spectral Signiture (SS) (Tran et al., 2018), STRIP (Gao et al., 2019), Anti-Backdoor Learning (ABL) (Li et al., 2021a) and Neural Cleanser (NC) (Wang et al., 2019). We follow the default settings for backbone attacks and defenses (see Appendix A.2). For CBS, we set $\epsilon=0.2$ and the corresponding poison rate is $0.2\%$ applied for Random and FUS. We retrain victim models on poisoned training data from scratch via SGD for 200 epochs with an initial learning rate of 0.1 and decay by 0.1 at epochs 100 and 150. Then we compare the success rate which is defined as the probability of classifying samples with triggers as the target class. We repeat every experiment 5 times and report average success rates (ASR) as well as the standard error if not specified. Results on Cifar10 are shown in Table 1 and results on Cifar100 are shown in Appendix A.4.
@@ 5.2 Performance of CBS in Type I backdoor attacks
Model ResNet18 ResNet18 normal-$\rightarrow$ VGG16
@@
Performance comparsion. Generally, CBS enhances the resilience of backbone attacks against various defense mechanisms. It achieves notable improvement compared to Random and FUS without a significant decrease in ASR when there are no defenses in place. This is consistent with our analysis in Section 4.3. We notice that though CBS has the lowest success rate when no defenses are active, CBS it still manages to achieve commendable performance, with success rates exceeding $70\%$ and even reaching $90\%$ for certain attacks. These indicate that CBS achieves a better trade-off between stealthiness and performance. It’s important to note that the effectiveness of CBS varies for different attacks and defenses. The improvements are more pronounced when dealing with stronger defenses and more vulnerable attacks. For instance, when facing SS, which is a robust defense strategy, CBS significantly enhances ASR for nearly all backbone attacks, especially for BadNet. In this case, CBS can achieve more than a $20\%$ increase compared to Random and a $15\%$ increase compared to FUS. Additionally, it’s worth mentioning that the first two defense mechanisms rely on detecting outliers in the latent space, and CBS consistently strengthens resistance against these defenses. This further supports the notion that boundary samples are inherently more challenging to detect and counteract. While the improvement of CBS on VGG16 is slightly less pronounced than on ResNet18, it still outperforms Random and FUS in nearly every experiment. This indicates that CBS can be effective even on unknown models.
5.3 Performance of CBS in Type II backdoor attacks
Attacks & Defenses. We consider 2 representative attacks in this category—Hidden-trigger (Saha et al., 2020) which adds imperceptible perturbations to samples to inject backdoors, and Clean-label (LC) (Turner et al., 2019) which leverages adversarial examples to train a backdoored model. We follow the default settings in the original papers, and adapt $l_{2}$-norm bounded perturbation (perturbation size $6/255$) for LC. We test all attacks against three representative defenses that are applicable to these attacks. We include Neural Cleanser (NC) (Wang et al., 2019), Spectral Signature (SS) (Tran et al., 2018), Fine Pruning (FP) (Liu et al., 2018), Anti-Backdoor Learning (ABL) (Li et al., 2021a). Details of these attacks and defenses are shown in Appendix A.2. We set $\epsilon=0.3$ for CBS and $p=0.2\%$ for Random and FUS correspondingly. For every experiment, a source class and a target class are randomly chosen, and poisoned samples are selected from the source class. The success rate is defined as the probability of misclassifying samples from the source class with triggers as the target class. Results on dataset Cifar10 and Cifar100 are presented in Table 2.

### v2

@@ 4.3 Theoretical understandings
To better understand CBS, we conduct theoretical analysis on a simple SVM model. As shown in Figure 3, in a 2-dimensional space, we consider a binary classification task where two classes are uniformly distributed in two balls centered at $\mu_{1}$(orange circle) and $\mu_{2}$(blue circle) with radius $r$ respectively in latent space [^6] This analysis is suitable for any neural networks whose last layer is a fully connected layer.:
$C_{1}\sim p_{1}(x)=\frac{1}{\pi r^{2}}1[\|x-\mu_{1}\|_{2}\leq r],\;\text{and }$ (3)
$C_{2}\sim p_{2}(x)=\frac{1}{\pi r^{2}}1[\|x-\mu_{2}\|_{2}\leq r],$
where let $\mu_{2}=0$ for simplicity. Assume that each class contains $n$ samples. We consider a simple attack that selects one single sample $x$ from class $C_{1}$, add a trigger to it to generate a poisoned $\tilde{x}$, and assign a label as class $C_{2}$ for it. Let $\tilde{C}_{1},\tilde{C}_{2}$ denote the poisoned data, and we can obtain a new backdoored decision boundary of SVM on the poisoned data. To study the backdoor effect of the trigger, we assume $\tilde{x}=x+\epsilon t/\|t\|$ where $t/\|t\|,\epsilon$ denote the direction and strength of the trigger, respectively.
@@
To explain this design, we assume that the trigger introduces a ’feature’ to the original samples (Khaddaj et al., 2023), and this ’feature’ is closely related to the target class while nearly orthogonal to the prediction features [^7] Prediction feature here is referred to features used for prediction when no triggers involved.. In addition, we assume $t$ is fixed for simplicity, which means this trigger is universal and we argue that this is valid because existing attacks such as BadNet (Gu et al., 2017) and Blend (Chen et al., 2017) inject the same trigger to every sample.To ensure the backdoor effect, we further assume $(\mu_{2}-\mu_{1})^{T}t\geq 0$, otherwise the poisoned sample will be even further from the target class (shown as the direction of the green dashed arrow) and lead to subtle backdoor effects. We are interested in two questions: (Q1) Are boundary samples harder to detect? (Q2) How do samples affect the backdoor performance?
@@
To investigate (Q1), we adopt the Mahalanobis distance (Mahalanobis, 2018) between the poisoned sample $\tilde{x}$ and the target class $\tilde{C}_{2}$ as an indicator of outliers. A smaller distance means $\tilde{x}$ is less likely to be an outlier, indicating better stealthiness. For (Q2), we estimate the success rate by estimating the volume (or area in 2D data) of the shifted class $C_{1}$ to the right of the backdoored decision boundary. This is because when triggers are added to every sample, the whole class will shift in the direction of $t$, shown as the orange dashed circle in Figure 3. The following series of theorems and propositions answer the above two questions. We begin with the Manahalnobis distance:
Theorem 4.2 (Mahalanobis distance).
@@
Assume $\tilde{x}=x+\epsilon t/\|t\|_{2}:=x+a$ for some arbitrary $x$ and some trigger $t$ and strength $\epsilon$. Also assume $\mu_{2}=0$, $(\mu_{2}-\mu_{1})^{T}\tilde{x}\geq 0$. Denote $\tilde{\mu}_{2}$ and $\tilde{S}_{2}$ as the sample mean and covariance matrix of the poisoned data with label $C_{2}$. Then the Mahalanobis distance between $\tilde{x}$ and the target class $\tilde{C}_{2}$ is defined as
$d_{M}^{2}(\tilde{x},\tilde{C}_{2})=(\tilde{x}-\tilde{\mu}_{2})^{T}\tilde{S}_{2}^{-1}(\tilde{x}-\tilde{\mu}_{2}).$
There exists some large constant $n_{0}$ so that when $n\geq n_{0}$, $d_{M}^{2}(\tilde{x},\tilde{C}_{2})$ satisfies
$P\left(\left|d_{M}^{2}(\tilde{x},\tilde{C}_{2})-\frac{4\|\tilde{x}\|^{2}}{r^{2}}\right|\geq t\right)\leq c_{1}\exp\left(-c_{2}t^{2}n\right)$ (4)
for some positive constants $c_{1}$ and $c_{2}$.
The proof of Theorem 4.2 can be found in Appendix 8.1. Since $\tilde{\mu}_{2}$ and $\tilde{S}_{2}$ are sample mean and covariance, we use concentration inequalities for vector average and matrix average to describe their behavior.
Theorem 4.2 shows the Mahalanobis distance ($d_{M}^{2}(\tilde{x},\tilde{C}_{2})$) from the fixed $\tilde{x}$ to $\tilde{C}_{2}$. When $n$ increases, $d_{M}^{2}(\tilde{x},\tilde{C}_{2})$ converges to its limit $4\|\tilde{x}\|^{2}/r^{2}$. In addition, to answer (Q1), the limit $4\|\tilde{x}\|^{2}/r^{2}$ is determined by the distance between $\tilde{x}$ and $\mu_{2}(=0)$: A larger $\|\tilde{x}\|$ results in a larger Mahalanobis distance, leading to a higher chance of $\tilde{x}$ being detected as an outlier (i.e., identified by the defender). To compare CBSand random selecting, if the attacker selects the support vector to attack, then $4\|\tilde{x}\|^{2}/r^{2}$ is minimized. Otherwise, $4\|\tilde{x}\|^{2}/r^{2}$ will be larger.
In the following, to answer (Q2), we first discuss the attack success rate assuming infinite training samples and then study how the finite sample affects the selection of the poisoned data and the margin of SVM.
Theorem 4.3 (Attack success rate, population).
Under the same conditions as Theorem 4.2, assume $n\rightarrow\infty$ and a hard margin exists, then the success rate for an arbitrary $\tilde{x}=x+a$ is an increasing function of
$\epsilon\cos(a,\tilde{x}-\mu_{1})-\|\tilde{x}-\mu_{1}\|/2-r/2.$
The proof of Theorem 4.3 can be found in Appendix 8.1. To figure out the attack success rate, we directly calculate the area of incorrect classification.
Remark 4.4 (Existence of hard margin).
Theorem 4.3 describes how the attack on $x$ affects the attack success rate. However, the existence of a hard margin depends on the selected sample $x$ and trigger $t$, which in turn determines whether an attack is effective. We provide Figure 4 for a better illustration. For the trigger $t$, we need $t^{T}(-\mu_{1})>0$, which indicates that the trigger $t$ moves the clean sample $x$ (from $C_{1}$) towards the target cluster $C_{2}$. To guarantee a hard margin, we require the poisoned sample $\widetilde{x}$ to fall into the shaded area, which is determined by $f_{1},f_{2}$ and $C_{1}$. $f_{1},f_{2}$ are the common tangent lines to $C_{1},C_{2}$, and are two extreme hyperplane that separates two clusters. Formally, we can define the area as $x\in\{x|\left(\{\|x+a-\mu_{1}\|_{2}\geq r\}\cap\{f_{1}(x+a)<0\}\cap\{f_{2}(x+a)>0\}\right)\cup\{f_{1}(x+a)>0\}\cup\{f_{2}(x+a)<0\}\}$. These conditions indicate that samples closer to the decision boundary are more likely to result in a hard margin in the poisoned data and guarantee an effective attack.
Figure 4: An illustrating figure for the existence of hard margin. The shaded area represents the region of $\tilde{x}$ where a hard margin exists.
On the other hand, unlike random sampling, since CBSrelies on the estimate of the confidence, we provide the following results to illustrate the impact of a finite sample size on CBSand random sampling. We first present Theorem 4.5 below to explain the change of the SVM margin under the finite-sample scenario using clean data:
Theorem 4.5 (Finite-sample scenario).
Under the same conditions as Theorem 4.2, consider the classification with clean data. Assume there are $n$ samples from $C_{1}$ and $n$ samples from $C_{2}$. Take $\delta_{n}=(\log n)/\sqrt{n}$. With probability at least
$1-2\left(1+2\left\lceil\frac{d_{1,f}-d_{1,\min}}{\delta_{n}}\right\rceil\right)\left(1-\frac{c\log^{2}n}{n\pi r^{2}}\right)^{n},$
uniformly for all hyperplane which separates $C_{1}$ and $C_{2}$, the corresponding margin for the $2n$ samples is $O(\sqrt{\delta_{n}})$-close to the margin for $C_{1}$ and $C_{2}$. The terms $d_{1,f}$ and $d_{1,\min}$ are constant values, and their definition can be found in (8) and (7) in Appendix 8.1.
The proof of Theorem 4.5 is in Appendix 8.1. The general idea is to construct some regions in $C_{1}$ and $C_{2}$ along their boundary and demonstrate that with a high probability, there is at least one sample that falls in each of the regions. Then we use these regions to quantify the difference between the margin to the population and the margin to the finite samples.
Based on Theorem 4.5, with a large enough $n$, the margin of SVM converges in probability. While Theorem 4.5 describes the margin using clean training data, the results can be further extended to discuss the sample selected by CBS, as well as the margin under poisoned data. In the following, Proposition 4.6 is for CBS, and Proposition 4.7 is for random sampling.
Proposition 4.6 (CBSin finite-sample scenario).
Under the conditions in Theorem 4.5, with the same probability as in Theorem 4.5, the sample selected by CBSis $O(\sqrt{\delta_{n}})$-close to the support vector in the population. If $t$ and $x$ are chosen such that the hard margin exists following Remark 4.4, the margin of the decision boundary determined by the finite poisoned samples is $O(\sqrt{\delta_{n}})$-close to its population version.
Proposition 4.7 (Random poisoning in finite-sample scenario).
Under the conditions in Theorem 4.5, with the same probability as in Theorem 4.5, if $t$ and $x$ are chosen such that the hard margin exists following Remark 4.4, then the margin of the decision boundary determined by the finite poisoned samples is $O(\sqrt{\delta_{n}})$-close to its population version.
Proposition 4.6 and 4.7 shows the consistency of these methods: When $n\rightarrow\infty$, the sample selected by CBSis close to the support vector in the population version. In addition, the margin of both methods converges to their population version respectively.
Remark 4.8 (Effectiveness-stealthiness trade-off).
@@
Based on the theorem, a smaller $\|\tilde{x}\|_{2}$ results in a smaller $d_{M}^{2}$, reducing the likelihood of being detected as an outlier. Additionally, closer proximity between $\tilde{x}$ and $\mu_{1}$ corresponds to a higher success rate without defenses. These observations highlight the trade-off between stealthiness and backdoor performance without defenses. Our experiments in Section 5 further demonstrate that incorporating boundary samples significantly improves stealthiness with only a slight reduction in success rate without defenses. [^8] Code can be found in https://github.com/PengfeiHePower/boundary-backdoor.
Remark 4.9 (Hard margin does not exist).
We also compare cases when the poisoned sample $x$ is too far from the target center $\mu_{2}$. When the poisoned sample is far enough from $\mu_{2}$ and the decision boundary, e.g., the poisoned sample $\widetilde{x}$, is still within reach of its true class, a hard margin will not exist. In this case, the misclassification of the single poisoned example will be ignored when $n$ is large enough, and the decision boundary of the SVM (with a soft margin) will be the same as the one from the clean SVM. Consequently, the poisoning effect is significantly reduced. To achieve a better success rate, the attacker needs to poison more samples, which can cause inefficiency and worse stealthiness. Therefore, poisoning samples closer to the boundary can even achieve better effectiveness while maintaining stealthiness.
@@ 5 Experiment
In this section, we conduct experiments to validate the effectiveness of CBS, and show its ability to boost the stealthiness of various existing attacks. We evaluate CBS and baseline samplings under no-defense and various representative defenses in Section 5.2 and 5.3. In particular, we select poisoned samples from the whole training data in these three sections and provide results when only partial data is accessible in Section 5.5 to validate the effectiveness of our approach in a broad and practical scenario. In Section 5.6, we will provide more empirical evidence to illustrate that CBS is harder to detect and mitigate. We also direct readers to additional experiments regarding larger datasets and more defenses in the Supplementary for a more comprehensive evaluation.
@@
Model ResNet18 ResNet18 $\rightarrow$ VGG16
@@
5.1 Experimental settings
To evaluate CBS and show its ability to be applied to various kinds of attacks, we consider 3 types  [^9] We determine the types based on the threat models of attacking methods. of attacking methods that cover most of existing backdoor attacks.
In detail, Type I backdoor attacks allow attackers to inject triggers into a proportion of training data and release the poisoned data to the public. Victims train models on them from scratch. The attack aims to misclassify samples with triggers as the pre-specified target class (also known as the all-to-one scenario). Type II backdoor attacks share the same threat model with Type I attacks, and the difference is that victims finetune pre-trained models on poisoned data, and the adversary’s goal is to misclassify samples from one specific class with triggers as the pre-specified target class (also known as the one-to-one scenario). Distinct from the preceding categories, Type III backdoor attacks necessitate an additional degree of control over the training process of the victim’s model. This control affords attackers the ability to concurrently optimize both the backdoor triggers and the model parameters, particularly in all-to-one attack scenarios.
Baselines for sampling. We compare CBS with two baselines—Random and FUS (Xia et al., 2022). The former selects samples to be poisoned with a uniform distribution, and the latter selects samples that contribute more to the backdoor injection via computing the forgetting events (Toneva et al., 2018) for each sample. In our evaluation, we focus on image classification tasks on datasets Cifar10 and Cifar100 (Krizhevsky et al., 2009) [^10] Additional datasets are included in Appendix 8.7, and model architectures ResNet18 (He et al., 2016), VGG16 (Simonyan & Zisserman, 2014). We use ResNet18 as the surrogate model 111111 Discussion of surrogate models in Appendix 8.10 for CBS and FUS if not specified. The surrogate model is trained on the clean training set via SGD for 60 epochs, with an initial learning rate of 0.01 and reduced by 0.1 after 30 and 50 epochs. We implement CBS according to Algorithm.1 and follow the original setting in (Xia et al., 2022) to implement FUS, i.e., 10 overall iterations and 60 epochs for updating the surrogate model in each iteration. After the generation of poisoned samples, we test the attacking performance on ResNet18 (the same architecture as the surrogate model) as well as transferring to another model architecture VGG16 (denoted as ResNet18 $\rightarrow$ VGG16 in tables 123).
5.2 Performance of CBS in Type I backdoor attacks
Attacks & Defenses. We consider 3 representative attacks in this category—BadNet (Gu et al., 2017) which attaches a small patch pattern as the trigger to samples to inject backdoors into neural networks; Blend (Chen et al., 2017) which applies the image blending to interpolate the trigger with samples; and Adaptive backdoor [^12] Both Adaptive-Blend and Adaptive-Patch are included (Qi et al., 2022) which introduces regularization samples to improve the stealthiness of backdoors, as backbone attacks. We include 4 representative defenses: Spectral Signiture (SS) (Tran et al., 2018) and STRIP (Gao et al., 2019) which are outlier-detection-based defenses, Anti-Backdoor Learning (ABL) (Li et al., 2021a) and Neural Cleanser (NC) (Wang et al., 2019) which are not detection-based defenses. We follow the default settings for backbone attacks and defenses. For CBS, we set $\epsilon=0.2$ and the corresponding poison rate is $0.2\%$ applied for Random and FUS, to guarantee that poisoning rates are the same for all sampling methods. We retrain victim models on poisoned training data from scratch via SGD for 200 epochs with an initial learning rate of 0.1 and decay by 0.1 at epochs 100 and 150. Then we compare the success rate which is defined as the probability of classifying samples with triggers as the target class. We repeat every experiment 5 times and report average success rates (ASR) as well as the standard error if not specified  [^13] Accuracy on the clean samples are included in Appendix 8.7. Results on Cifar10 are shown in Table 1, and results on Cifar100 and Tiny-ImageNet are shown in the Appendix.